# LLM-Driven Multi-step Translation from C to Rust using Static Analysis

### v2

@@ Abstract
Translating software written in legacy languages to modern languages, such as C to Rust, has significant benefits in improving memory safety while maintaining high performance. However, manual translation is cumbersome, error-prone, and produces unidiomatic code. Large language models (LLMs) have demonstrated promise in producing idiomatic translations, but offer no correctness guarantees as they lack the ability to capture all the semantics differences between the source and target languages. To resolve this issue, we propose SACTOR, an LLM-driven C-to-Rust zero-shot translation tool using a two-step translation methodology: an “unidiomatic” step to translate C into Rust while preserving semantics, and an “idiomatic” step to refine the code to follow Rust’s semantic standards. SACTOR utilizes information provided by static analysis of the source C program to address challenges such as pointer semantics and dependency resolution. To validate the correctness of the translated result from each step, we use end-to-end testing via the foreign function interface to embed our translated code segment into the original code. We evaluate the translation of $200$ programs from two datasets and two case studies, comparing the performance of GPT-4o, Claude 3.5 Sonnet, Gemini 2.0 Flash, Llama 3.3 70B and DeepSeek-R1 in SACTOR. Our results demonstrate that SACTOR achieves high correctness and improved idiomaticity, with the best-performing model (DeepSeek-R1) reaching 93% and (GPT-4o, Claude 3.5, DeepSeek-R1) reaching 84% correctness (on each dataset, respectively), while producing more natural and Rust-compliant translations compared to existing methods.
@@ 1 Introduction
The C language is the most widely used language in system-level programming. C’s ability to directly manipulate memory and hardware has made it indispensable in developing operating systems, embedded systems, and performance-critical applications (Love, 2013). However, its manual memory management has led to memory-related bugs, such as buffer overflows, dangling pointers, and memory leaks, which have been the root cause of many security vulnerabilities and system crashes (Fan et al., 2020).
To address these shortcomings, Rust has emerged as a promising alternative to C. Rust is a modern system programming language that offers memory safety without relying on garbage collection by imposing a strict object ownership model (Matsakis and Klock, 2014). Due to Rust’s memory safety guarantees and performance, it is adopted by projects like the Linux kernel in developing new hardware drivers [^1] https://github.com/Rust-for-Linux/linux and Mozilla Firefox. Translating legacy C code to Rust provides a path to improve safety and performance, but manual translation is time-intensive, error-prone, and requires developer expertise in both languages.
Researchers have started investigating automatic translation methods for C-to-Rust, resulting in tools such as C2Rust (Immunant, 2020) which analyzes the abstract syntax tree (AST) of the code to produce syntactically equivalent Rust code. Such “traditional” approaches for C-to-Rust translation often rely on rule-based techniques or static analysis (Zhang et al., 2023; Immunant, 2020; Emre et al., 2021; Hong and Ryu, 2024; Ling et al., 2022). While these methods are effective at producing syntactically correct translations, they often fail to provide an “idiomatic” translation: one that follows the Rust programming model and conventions with the least use of unsafe blocks. Due to fundamental differences in semantics and memory management between C and Rust, idiomatic translations are crucial to compiler-enforced memory safety, as well as improving code readability and maintainability.
Large language models (LLMs) have demonstrated the ability to understand syntax and code semantics, making them a promising tool for code translation (Pan et al., 2023). However, LLMs are prone to hallucinations and often generate incorrect code (Perry et al., 2023). Moreover, the large semantic differences between C and Rust make it more challenging for LLMs to capture correct semantics of the code, leading to inaccurate translations. Previous work has explored various verification techniques, such as fuzz tests and symbolic execution (Yang et al., 2024; Eniser et al., 2024) to validate LLM-generated code. Although these methods can guarantee equivalence, their strictness often limits scalability or applicability to code with complicated code features. For example, Vert (Yang et al., 2024) always struggles to verify the correctness of code with data structures through symbolic execution.
In this paper, we introduce Structure-Aware C-to-Rust Translator (or SACTOR), an LLM-driven zero-shot translation tool. An overview of SACTOR is presented in Figure 2. We divide the translation process into two steps: 1. C to unidiomatic Rust where we focus on syntactic translation only and allow C-like semantics and unsafe code blocks to keep all the unsafe memory manipulations, and 2. unidiomatic Rust to idiomatic Rust where we refine the translation to follow Rust semantic standards and remove all the unsafe blocks in the code to ensure safety. In both translation steps, we leverage static analysis tools to provide critical information on pointers and code dependencies and guide the translation process. To address the challenge of verifying LLM outputs efficiently, we utilize end-to-end testing by embedding the translated Rust code with the original C code through foreign function interface (FFI). This allows us to directly run the end-to-end tests on the whole program to validate a particular translation. This combination of static analysis, multi-step refinement, and efficient (and robust) verification significantly improves translation quality, achieves high-level correctness, and provides a broader applicability to different types of C programs.
We make the following contributions:
  • We propose SACTOR, a tool that integrates feedback from static analysis tools into a two-step translation process. By doing so, we reliably obtain idiomatic unsafe-free translated Rust from C (§ 4).
  • We propose and demonstrate a new verification approach by embedding the translated code into the original code and utilizing the end-to-end testing. This enables more practical, scalable, and efficient verification method for LLM-generated code (§ 4.3).
  • We evaluate our approach on a diverse set of C programs across two datasets and five different LLMs. Our best-performing model, DeepSeek-R1 achieves 93% success rate on the first dataset, while GPT-4o, Claude 3.5, and DeepSeek-R1 reach 84% on the second dataset (§ 6.1.1). Notably, SACTOR generates more idiomatic, unsafe-free Rust code compared to four existing methods (§ 6.3). These results demonstrate the broad applicability and effectiveness of SACTOR. [^2] SACTOR code is available at https://github.com/qsdrqs/sactor and datasets are available at https://github.com/qsdrqs/sactor-datasets.
  • We compare the cost of each LLM in terms of token consumption and average queries required for a single translation and observe that GPT-4o is the most efficient in terms of token consumption while DeepSeek-R1 consumes 5-7$\times$ more tokens than GPT-4o. On the other hand, We observe that the average number of queries varies across models but is less pronounced than token consumption. The most query-intensive model, Llama 3.3, requires only about 1.4$\times$ more queries than the least costly model, Gemini 2.0 (§ 6.2).
  • Finally, we inspect failure cases and find that different LLMs exhibit varying strengths and weaknesses in translating C code to Rust (§ 6.1.2). DeepSeek-R1 excels in handling format string errors, array manipulation, and has a lower failure rate in different categories of errors on the TransCoder-IR dataset, which features complex C code. This suggests reasoning models like DeepSeek-R1 perform better in translating code with intricate logic and data manipulations.

### v3

# SACTOR: LLM-Driven Correct and Idiomatic C to Rust Translation with Static Analysis and FFI-Based Verification

@@ Abstract
[…]
@@ 1 Introduction
C is widely used due to its ability to directly manipulate memory and hardware (love2013linux). However, manual memory management leads to vulnerabilities such as buffer overflows, dangling pointers, and memory leaks (bigvul). Rust addresses these issues by enforcing memory safety through a strict ownership model without garbage collection (matsakis2014rust), and has been adopted in projects like the Linux kernel [^1] https://github.com/Rust-for-Linux/linux and Mozilla Firefox. Translating legacy C code into idiomatic Rust improves safety and maintainability, but manual translation is error-prone, slow, and requires expertise in both languages.
Automatic tools such as C2Rust (c2rust) generate Rust by analyzing C ASTs, but rule-based or static approaches (crown; c2rust; emre2021translating; hong2024don; ling2022rust) typically yield unidiomatic code with heavy use of unsafe. Given semantic differences between C and Rust, idiomatic translations are crucial for compiler-enforced safety, readability, and maintainability.
Large language models (LLMs) show potential for capturing syntax and semantics (pan2023understanding), but they hallucinate and often generate incorrect or unsafe code (perry2023users). In C-to-Rust translation, naive prompting produces unsafe or semantically misaligned outputs. Prior work has explored prompting strategies (syzygy; c2saferrust; shiraishi2024context) and verification methods such as fuzzing and symbolic execution (vert; flourine). While these improve correctness, they struggle with complex programs and rarely yield idiomatic Rust. For example, Vert (vert) fails on programs with complex data structures, and C2SaferRust (c2saferrust) still produces Rust with numerous unsafe blocks.
In this paper, we introduce SACTOR, a structure-aware, LLM-driven C-to-Rust translator (Figure 1). SACTOR follows a two-stage pipeline:
  • 1.
    C $\to$ Unidiomatic Rust: Interface-preserving translation that may use unsafe for low-level operations.
  • 2.
    Unidiomatic $\to$ Idiomatic Rust: Behaviorally-equivalent translation that refines to Rust idioms, eliminating unsafe and migrating C API patterns to Rust equivalents.
Static analysis of C code (pointer semantics, dependencies) guides both stages. To verify correctness, we embed the translated Rust with the original C via the Foreign Function Interface (FFI), enabling end-to-end testing on both stages and accept a stage when all end-to-end tests can pass. This decomposition separates syntax from semantics, simplifies the LLM task, and ensures more idiomatic, memory-safe Rust [^2] SACTOR code is available at https://github.com/qsdrqs/sactor and datasets are available at https://github.com/qsdrqs/sactor-datasets. An example of SACTOR translation process is in Appendix E.
LLM orchestration. SACTOR places the LLM inside a neuro-symbolic feedback loop. Static analysis and a machine-readable interface specification guide prompting; compiler diagnostics and end-to-end tests provide structured feedback. In the idiomatic verification phase, a rule-based harness generator with an LLM fallback completes the feedback loop. This design first ensures semantic correctness in unidiomatic Rust, then refines it into idiomatic Rust, with both stages verifiable in a unified two-step process.
Our contributions are as follows:
  • Method: An LLM-orchestrated, structure-aware two-phase pipeline that separates semantic preservation from idiomatic refinement, guided by static analysis (§ 4)
  • Verification: SACTOR verifies both unidiomatic and idiomatic translations via FFI-based testing. During idiomatic verification, it uses a co-produced interface specification to synthesize C/Rust harnesses with an LLM fallback for missing patterns; compiler and test feedback are structured into targeted prompt repairs (§ 4.3).
  • Evaluation: Across two datasets (200 programs) and five LLMs, SACTOR reaches 93% / 84% end-to-end correctness (DeepSeek-R1) and improves idiomaticity (§ 6.2). On CRust-Bench (50 samples), unidiomatic translation averages 85% function-level success rate across all samples (82% aggregated across functions), with 32/50 samples fully translated; idiomatic success is computed on those 32 samples and averages 52% (43% aggregated; 8/32 fully idiomatic). On libogg (77 functions), the function-level success rate is 100% for unidiomatic and 53% and 78% for idiomatic across GPT-4o and GPT-5, respectively (§ 6.3).
  • Diagnostics: We analyze efficiency, feedback, temperature sensitivity, and failure cases: GPT-4o is the most token-efficient, compilation/testing feedback boosts weaker models by 17%, temperature has little effect, and reasoning models like DeepSeek-R1 excel on complex bugs such as format-string and array errors (Appendix H).
Figure 1: Overview of the SACTOR methodology.